What are the Harley Davidson fixed asset inventory turnover ratio, fixed asset turnover ratio, total asset turnover ratio, debt ratio, equity multiplier ratio, times interest earned ratio, profit margin ratio,return on assets ratio, return on equity ratio, price earnings ratio, current ratio, quick ratio,  for 2016, 2017, and 2018? I do not know how to pull the information from the 10k as the terminology on the 10k is different than the formula terminology.
Answer #1

We use the publicly available financials of Harley Davidson for the years 2018, 2017 and 2016 to calculate the ratios :

Inventory turnover ratio = cost of revenue / inventory

2018 = 3,803,576 / 556,128 = 6.84

2017 = 3,739,199 / 538,202 = 6.95

2016 = 3,885,527 / 499,917 = 7.77

Fixed asset turnover ratio = revenue / fixed assets

2018 = 5,716,875 / 904,132 = 6.32

2017 = 5,647,224 / 967,781 = 5.84

2016 = 5,996,458 / 981,593 = 6.11

total asset turnover ratio = revenue / total assets

2018 = 5,716,875 / 10,665,664 = 0.54

2017 = 5,647,224 / 9,972,672 = 0.57

2016 = 5,996,458 / 9,890,240 = 0.61

debt ratio = total liabilities / total assets

2018 = 8,891,715 / 10,665,664 = 0.83

2017 = 8,128,395 / 9,972,672 = 0.82

2016 = 7,970,082 / 9,890,240 = 0.81